# VDB: A Deep-Learning Framework for Sparse, Large-Scale, and High-Performance Spatial Intelligence

### v1

@@ Abstract.
We present VDB, a novel GPU-optimized framework for deep learning on large-scale 3D data. VDBÂ provides a complete set of differentiable primitives to build deep learning architectures for common tasks in 3D learning such as convolution, pooling, attention, ray-tracing, meshing, etc. VDBÂ simultaneously provides a much larger feature set (primitives and operators) than established frameworks with no loss in efficiency: our operators match or exceed the performance of other frameworks with narrower scope. Furthermore, VDBÂ can process datasets with much larger footprint and spatial resolution than prior works, while providing a competitive memory footprint on small inputs. To achieve this combination of versatility and performance, VDBÂ relies on a single novel VDB index grid acceleration structure paired with several key innovations including GPU accelerated sparse grid construction, convolution using tensorcores, fast ray tracing kernels using a Hierarchical Digital Differential Analyzer algorithm (HDDA), and jagged tensors. Our framework is fully integrated with PyTorch enabling interoperability with existing pipelines, and we demonstrate its effectiveness on a number of representative tasks such as large-scale point-cloud segmentation, high resolution 3D generative modeling, unbounded scale Neural Radiance Fields, and large-scale point cloud reconstruction.
@@ 3.5. Accelerated Sparse Convolutional Operators
VDBÂ has been designed to be compatible with highly efficient algorithms for convolutional operations on sparse data, such as the Sorted Implicit Gemm (SpConv v2) paradigm used in TorchSparse++. We emphasize that leveraging such highly-tuned libraries in the context of our hierarchical, tree-based indexing structure is a straightforward exercise: VDBÂ is effectively a locality-optimizing mapping between a sparse collection of lattice indices and a one-dimensional, linear index space. Contrary to random hash-based maps, VDBÂ inherently provides the property that active indices that are geometrically proximate in the containing 3D lattice, will have high probability of also being proximate in linear index space. Conversely, active voxels corresponding to a contiguous sub-sequence of linear indices are highly likely to be geometrically clustered together in the containing 3D lattice. Other than this (favorable) inherent property of the VDBÂ indexing scheme, our data structure is drop-in compatible with implementations that originate from hash-based structures (e.g. SpConv v2) by simply treating the linear index of each active voxel as a âhash keyâ (but with built-in locality properties). We have incorporated SpConv v2 into our operator toolkit and, as our micro-benchmarks reveal, we at minimum match the efficiency of TorchSparse++ at the operator level within our framework.
@@
In light of this, we consider an alternative where we prioritize regularity over sparsity of computation, essentially tolerating a higher compute burden for the sake of more local structure. Specifically, we have implemented a kernel that performs local densification in GPU shared memory, at the level of an VDBÂ leaf node, and performs a fully regular and (locally) dense convolution within this window. In detail, we allocate space in shared memory for a locally densified copy of the input activations in a window of size stradding the leaf node, plus a one-voxel halo in its immediate neighborhood (a footprint of 31.25KB for 4-byte FP32/TF32 data, at feature width of 8). Likewise, the output of this operation is an buffer of 16-wide output feature vectors (footprint of 32KB) also stored in shared memory. We subdivide the local domain into 32 subtiles, assign each of them to a warp (1024 total threads) and use WMMA tensorcore GEMMs (at TF32 precision with FP32 accumulate) within each warp to apply each of the 27 spokes of the stencil. Even though this paradigm clearly performs more computation than strictly necessary (foregoing sparity due to either voxel or stencil occupancy), the regularity of the computation in combination with the memory-bound nature of this scenario allows for superior performance (relative to our SpConv v2 default backend) in leaf nodes that have an occupancy of 20% or higher (all the way to an approximately 2.5x-3x advantage for a dense domain). It should also be noted that no auxiliary indexing structures are necessary for this kernel approach, all gather offsets are computed directly and efficiency from the (very lightweight) metadata of the core VDBÂ tree structure, taking advantage of amortization. Finally, due to the compact and local storage of all (output) feature vectors within a leaf node, the writeback of the convolution result into global memory occurs on a fully sequential memory range (all active indices within a leaf node are sequentially indexed).
@@
(a) We block the gather-GEMM-scatter operation so that it is performed on a contiguous subsequence of output indices from the VDBÂ data structure, typically 64 indices at a time. Due to the locality of the VDBÂ mapping, those indices are expected to correspond to highly clustered geometric coordinates from one or more IndexGrid leaf nodes.
@@ 3.6.2. List of Operators
A common operator is to sample tensor values on a voxel grid at a set of query points . VDB provides differentiable sampling operators which accept a GridBatch , a JaggedTensor of per-voxel features with shape , and a JaggedTensor of query points with shape . These operators return a set of features sampled at each point using Trilinear or BÊzier interpolation.
@@
VDB supports splatting data stored at points onto a grid using Trilinear or BÊzier interpolation. These operators accept a GridBatch , a JaggedTensor of points, and a JaggedTensor of per-point features. They produce a JaggedTensor of features (one per voxel in ) by splatting the feature at each point onto the neighboring voxels.
@@ 4. Experiments
In this section, we demonstrate the effectiveness of VDBÂ through a series of benchmarks and qualitative examples of use cases. Our experiments demonstrate that our framework successfully covers a broad variety of use cases and operations, while achieving state-of-the-art runtime performance and memory efficiency. First, we perform micro-benchmarks of the most important operators in VDB, comparing them against corresponding state-of-the-art operators in other sparse deep learning frameworks in terms of both memory usage and speed. Next, we run a macro-benchmark showing that VDBÂ remains performant in the real-world use case of training a sparse convolutional neural network (CNN). Finally, we demonstrate the utility of VDBÂ by showing its use in several key applications on high-resolution 3D data. These applications include 3D reconstruction from points, semantic completion, 3D shape generation, and neural radiance field rendering.
@@ 4.1.1. IndexGrid Construction
The IndexGrid construction algorithm, detailed in Section 3.3, converts a list of ijk integer or xyz point coordinates into a VDB IndexGrid on the GPU. It forms the backbone of all grid constructions in VDB, while also acting as a means to initialize sparse grids. We evaluate the runtime performance and memory footprint of our grid construction algorithm against those in TorchSparse++ (Tang et al., 2023), Minkowski Engine (Choy et al., 2019), and spconv (Contributors, 2022) by constructing a grid with random points sampled from a normal distribution. Figure 9 shows the maximum memory usage and runtime when constructing a grid from an increasing number of input points. Our method is comparable to baselines in terms of runtime performance while offering significant advantages in terms of memory efficiency. We remark that the three baseline approaches run out of memory long before ours. Thus, VDB can process much larger input data than current state-of-the-art sparse DL frameworks.
@@ 5. Conclusion and Future Work
We presented VDB, a novel GPU-optimized framework for deep learning on large-scale 3D data. Our framework includes a broad set of novel differential primitives which can be used to build deep-learning pipelines for a wide variety of 3D tasks. These primitives include GPU accelerated grid building, ray marching, convolution, sampling, splatting, etc. Furthermore, VDBÂ has a significantly more comprehensive suite of features than existing frameworks, runtime performance that is at-par or superior to state-of-the-art and memory efficiency that exceeds state-of-the-art by a large margin. VDBÂ uses a single, novel VDB IndexGrid data structure to accelerate all operations, making it composable and easily extensible. We demonstrated the effectiveness of VDBÂ via extensive quantitative benchmarks and qualitative demonstrations on real-world 3D learning use cases, showing that VDBÂ enables high-performance deep learning on large scale 3D data.

### v2

@@ Abstract.
We present VDB, a novel GPU-optimized framework for deep learning on large-scale 3D data. VDB provides a complete set of differentiable primitives to build deep learning architectures for common tasks in 3D learning such as convolution, pooling, attention, ray-tracing, meshing, etc. VDB simultaneously provides a much larger feature set (primitives and operators) than established frameworks with no loss in efficiency: our operators match or exceed the performance of other frameworks with narrower scope. Furthermore, VDB can process datasets with much larger footprint and spatial resolution than prior works, while providing a competitive memory footprint on small inputs. To achieve this combination of versatility and performance, VDB relies on a single novel VDB index grid acceleration structure paired with several key innovations including GPU accelerated sparse grid construction, convolution using tensorcores, fast ray tracing kernels using a Hierarchical Digital Differential Analyzer algorithm (HDDA), and jagged tensors. Our framework is fully integrated with PyTorch enabling interoperability with existing pipelines, and we demonstrate its effectiveness on a number of representative tasks such as large-scale point-cloud segmentation, high resolution 3D generative modeling, unbounded scale Neural Radiance Fields, and large-scale point cloud reconstruction.
@@ 3.5. Accelerated Sparse Convolutional Operators
VDB has been designed to be compatible with highly efficient algorithms for convolutional operations on sparse data, such as the Sorted Implicit Gemm (SpConv v2) paradigm used in TorchSparse++. We emphasize that leveraging such highly-tuned libraries in the context of our hierarchical, tree-based indexing structure is a straightforward exercise: VDB is effectively a locality-optimizing mapping between a sparse collection of lattice indices and a one-dimensional, linear index space. Contrary to random hash-based maps, VDB inherently provides the property that active indices that are geometrically proximate in the containing 3D lattice, will have high probability of also being proximate in linear index space. Conversely, active voxels corresponding to a contiguous sub-sequence of linear indices are highly likely to be geometrically clustered together in the containing 3D lattice. Other than this (favorable) inherent property of the VDB indexing scheme, our data structure is drop-in compatible with implementations that originate from hash-based structures (e.g. SpConv v2) by simply treating the linear index of each active voxel as a âhash keyâ (but with built-in locality properties). We have incorporated SpConv v2 into our operator toolkit and, as our micro-benchmarks reveal, we at minimum match the efficiency of TorchSparse++ at the operator level within our framework.
@@
In light of this, we consider an alternative where we prioritize regularity over sparsity of computation, essentially tolerating a higher compute burden for the sake of more local structure. Specifically, we have implemented a kernel that performs local densification in GPU shared memory, at the level of an VDB leaf node, and performs a fully regular and (locally) dense convolution within this window. In detail, we allocate space in shared memory for a locally densified copy of the input activations in a window of size stradding the leaf node, plus a one-voxel halo in its immediate neighborhood (a footprint of 31.25KB for 4-byte FP32/TF32 data, at feature width of 8). Likewise, the output of this operation is an buffer of 16-wide output feature vectors (footprint of 32KB) also stored in shared memory. We subdivide the local domain into 32 subtiles, assign each of them to a warp (1024 total threads) and use WMMA tensorcore GEMMs (at TF32 precision with FP32 accumulate) within each warp to apply each of the 27 spokes of the stencil. Even though this paradigm clearly performs more computation than strictly necessary (foregoing sparity due to either voxel or stencil occupancy), the regularity of the computation in combination with the memory-bound nature of this scenario allows for superior performance (relative to our SpConv v2 default backend) in leaf nodes that have an occupancy of 20% or higher (all the way to an approximately 2.5x-3x advantage for a dense domain). It should also be noted that no auxiliary indexing structures are necessary for this kernel approach, all gather offsets are computed directly and efficiency from the (very lightweight) metadata of the core VDB tree structure, taking advantage of amortization. Finally, due to the compact and local storage of all (output) feature vectors within a leaf node, the writeback of the convolution result into global memory occurs on a fully sequential memory range (all active indices within a leaf node are sequentially indexed).
@@
(a) We block the gather-GEMM-scatter operation so that it is performed on a contiguous subsequence of output indices from the VDB data structure, typically 64 indices at a time. Due to the locality of the VDB mapping, those indices are expected to correspond to highly clustered geometric coordinates from one or more IndexGrid leaf nodes.
@@ 3.6.2. List of Operators
A common operator is to sample tensor values on a voxel grid at a set of query points . VDB provides differentiable sampling operators which accept a GridBatch , a JaggedTensor of per-voxel features with shape , and a JaggedTensor of query points with shape . These operators return a set of features sampled at each point using Trilinear or BĂŠzier interpolation.
@@
VDB supports splatting data stored at points onto a grid using Trilinear or BĂŠzier interpolation. These operators accept a GridBatch , a JaggedTensor of points, and a JaggedTensor of per-point features. They produce a JaggedTensor of features (one per voxel in ) by splatting the feature at each point onto the neighboring voxels.
@@ 4. Experiments
In this section, we demonstrate the effectiveness of VDB through a series of benchmarks and qualitative examples of use cases. Our experiments demonstrate that our framework successfully covers a broad variety of use cases and operations, while achieving state-of-the-art runtime performance and memory efficiency. First, we perform micro-benchmarks of the most important operators in VDB, comparing them against corresponding state-of-the-art operators in other sparse deep learning frameworks in terms of both memory usage and speed. Next, we run a macro-benchmark showing that VDB remains performant in the real-world use case of training a sparse convolutional neural network (CNN). Finally, we demonstrate the utility of VDB by showing its use in several key applications on high-resolution 3D data. These applications include 3D reconstruction from points, semantic completion, 3D shape generation, and neural radiance field rendering.
@@ 4.1.1. IndexGrid Construction
The IndexGrid construction algorithm, detailed in Section 3.3, converts a list of ijk integer or xyz point coordinates into a VDB IndexGrid on the GPU. It forms the backbone of all grid constructions in VDB, while also acting as a means to initialize sparse grids. We evaluate the runtime performance and memory footprint of our grid construction algorithm against the analogous calculations in TorchSparse++ (Tang et al., 2023), MinkowskiEngine (Choy et al., 2019), and spconv (Contributors, 2022) by constructing a grid with random points sampled from a normal distribution followed by any auxiliary computation required to execute the sparse convolution. Figure 9 shows the maximum memory usage and runtime of this process as a function of the number of input points. Our method runs significantly faster than MinkowskiEngine and spcov and on par with TorchSparse++. Moreover, our approach is more memory efficient than TorchSparse++ while MinkowskiEngine and spconv both fail as the problem size increases. Thus, VDB can process input data much larger than current state-of-the-art sparse DL frameworks. We conducted this experiment on an NVIDIA A100 GPU.
@@ 5. Conclusion and Future Work
We presented VDB, a novel GPU-optimized framework for deep learning on large-scale 3D data. Our framework includes a broad set of novel differential primitives which can be used to build deep-learning pipelines for a wide variety of 3D tasks. These primitives include GPU accelerated grid building, ray marching, convolution, sampling, splatting, etc. Furthermore, VDB has a significantly more comprehensive suite of features than existing frameworks, runtime performance that is at-par or superior to state-of-the-art and memory efficiency that exceeds state-of-the-art by a large margin. VDB uses a single, novel VDB IndexGrid data structure to accelerate all operations, making it composable and easily extensible. We demonstrated the effectiveness of VDB via extensive quantitative benchmarks and qualitative demonstrations on real-world 3D learning use cases, showing that VDB enables high-performance deep learning on large scale 3D data.